Let f(x) = 3x + 2 and g(x) 3x + 2 and g(x) = 4x2 + 2x. After simplifying, (fog)(x) =
Answer #1

(fog)(x) = Greiven that f(2)= 3x+2, 96) = 4x² + 2x f (glas) -#(4x2+2a) =3(4x²+2x)+2 = 1222 +6842. :foqxo3 = 124 62+2.